Roofers need to move equipment onto the roof of a building. They have a machine that will pull a platform at varying accelerations. The rope used to pull the platform will snap if a force of 5550 Newtons is exceeded. The platform itself has a mass of 50.0kg. The equipment has a mass of 475 kg. What is the maximum acceleration the machine could lift the equipment without snapping the rope?
1
Expert's answer
2020-11-16T07:53:24-0500

In order to prevent the snap, the acceleration should be (accroding to the Newton's second law):


"a\le\dfrac{F}{m}"

where "F = 5550N" is the force threshold, and "m = 50kg + 475kg = 525kg" is the total mass of the platform and equipment. Thus, obtain:


"a\le\dfrac{5550N}{525} = 10.6m/s^2"

Answer. The maximum acceleration is 10.6 m*s^-2.